# Optimal CO2 storage management considering safety constraints in multi-stakeholder multi-site CCS projects: a game theoretic perspective

### v1

@@ 4.3 Reward and Penalty Modeling for Geological CO2 storage
To address these challenges, we adopt coalition structures inspired by game theory, enabling flexible combinations of cooperative and competitive behaviors. Coalition structures are enumerated using Bell numbers, allowing evaluation of all possible collaborative configurations. In our case study with three operators (A, B, and C), two primary coalition structure are considered: fully cooperative where all companies act collectively to maximize the total net present value (NPV) of the project, and fully competitive where each company independently seeks to maximize its own NPV without regard for the others. This framework allows us to capture a wide spectrum of strategic behaviors, from joint optimization for mutual benefit to self-interested competition. Based on these coalition types, distinct reward and penalty structures are defined to reflect the corresponding operational incentives and constraints.:
  • 1.
    {A} v.s {B} v.s {C} setting: Each agent $i$ maximizes its own net present value ($NPV_{i}$), which is cumulative reward over the whole injection period, defined as:
    $NPV_{i}=\sum_{t=1}^{T}\gamma^{t}\cdot PV_{i}(t)=\sum_{t=1}^{T}\gamma^{t}\cdot r_{i}(t),$ (16)
    where $PV_{i}(t)$ accounts for tax credits and operational costs associated with CO2 storage and brine handling.
  • 2.
    {A & B & C} setting: Agents A, B and C form coalitions to maximize the team reward, defined as the sum of NPVs across coalition members.
    $NPV_{i}=\sum_{i\in\{A,B,C\}}\sum_{t=1}^{T}\gamma^{t}\cdot r_{i}(t),$ (17)
    This incentivizes coordinated well controls and pressure management strategies that improve collective outcomes while maintaining safety.
In the meanwhile, we consider two structures for the penalty model, the first structure only considers imposing penalty when the pore pressure of the injection well grid block exceed the threshold pressure
$c_{i}(t)=5000\cdot\sum_{g\in\mathcal{G}_{well,i}}\mathbf{1}\!\left(p_{g}^{t}>p_{threshold}\right),$ (18)
while the second structure considers a region-wide penalty, defined as:
$c_{i}(t)=50\cdot\sum_{g\in\mathcal{G}_{i}}\mathbf{1}\!\left(p_{g}^{t}>p_{threshold}\right),$ (19)
@@
In this formulation, $c_{i}^{t}$ denotes the penalty incurred by agent $i$ at time step $t$. The set $\mathcal{G}_{i}$ represents all grid blocks within company $i$’s project area, and $p_{g}^{t}$ is the pressure in grid block $g$ at time $t$. The indicator function $\mathbf{1}(p_{g}^{t}>=p_{i,\text{th}})$ returns $1$ if the pressure in grid block $g$ exceeds the threshold pressure $p_{i,\text{th}}$ assigned to agent $i$, and 0 otherwise. Thus, the summation counts the total number of violating grid blocks, which is then multiplied by a fixed penalty factor of $50$ (50 is selected in a way that the penalty value is of same magnitude of the reward in order for good training practice) to determine the agent’s cost at that time step.
@@
$\sum_{t=1}^{T}\gamma^{t}\cdot c_{i}(t)\leq d_{i}$
$d_{i}$ is the penalty threshold, in this experiment, because the pressure constraints are strictly imposed, $d_{i}$ is set to be 0.0.
@@ 5.1 Description of Reservoir and Model Setup
The proposed framework is implemented in a synthetic reservoir where the dimension spans $44\mathrm{km}\times 12.8\mathrm{km}\times 0.2\mathrm{km}$ in x, y and z directions. With a depth of 4,000m4{,}000\ $\mathrm${m} and an assumed fracture gradient of 17.0kPa/m17.0\ $\mathrm${kPa/m}, the fracture pressure $p_{\mathrm{frac}}$ is set to be 68,000kPa68{,}000\ $\mathrm${kPa}. For safe CO2 injection, the pore pressure must remain below the fracturing pressure, expressed as:
@@
where $0.0<\alpha<1.0$ reflects site-specific risk tolerance. In this study, $\alpha$ also accounts for the $\sim 3\%$ prediction error between the reduced-order proxy model and the high-fidelity model. The permeability distribution for the basin is shown in Figure 6.
We consider three companies (A, B, and C) operating in non-overlapping regions of the basin (as indicated in the shaded area in 6), each managing a different number of injection wells. To reflect varying risk tolerances, the maximum allowable pore pressure for each company is set to:
$p_{\mathrm{threshold},i}=\alpha_{i}\,p_{\mathrm{frac}}=\begin{cases}65{,}000\ \mathrm{kPa},&i=A,\\ 55{,}000\ \mathrm{kPa},&i=B,\\ 55{,}000\ \mathrm{kPa},&i=C.\end{cases}$
@@
Economic assumptions for revenue and cost calculations are summarized in Table 1, with all companies sharing the same unit revenue and cost structure.
@@
Figure 7 displays well control strategies of all 3 companies for 25 years, while figure 8 presents the corresponding reservoir pressure distribution at the end of injection. In the latter, annotations highlight the maximum grid-block pressures within each company’s leased area at selected injection stages. The results clearly show that under discretionary well control, all three companies operate beyond their allowable pressure thresholds, increasing the risk of caprock fracturing. This exceedance suggests that without coordinated control or imposing safety constraints explicitly, CCS operations may compromise long-term reservoir integrity and storage security.
@@
5.2 Cooperative vs. Competitive from MADDPG
In this subsection, we present the results for scenarios where companies either form a grand coalition to maximize total team NPV (cooperative case) or operate independently without any coalition (competitive case). Figure 9 illustrates the reward and penalty trajectories during training. Initially, both coalition structures adopt aggressive injection strategies, leading to penalties that exceed the predefined safety thresholds. Over time, the penalties decrease and remain below the threshold, indicating compliance with safety constraints. The intermittent pattern of the penalty curve arises from the sparse penalty formulation described in equation 18. While MADDPG does not guarantee a strictly Pareto-optimal policy, we identify the optimal policy as the one achieving the highest team reward while maintaining penalties below the threshold. Figure 10 summarizes the resulting NPVs for both coalition structures, showing that the total team NPV increases from \$6875.34 million in the competitive setting to \$8702.41 million in the cooperative setting. At the company level, company A and C gain substantial benefits from coalition formation, with NPV increases of \$302.33 million and \$1668.91 million, or 14.1% and 95.9% in terms of percentage respectively, compared to acting alone. In contrast, company B experiences a marginal decrease in NPV. Nevertheless, it would still have an incentive to participate in the coalition if an equitable profit-sharing mechanism were implemented, ensuring mutual benefits across all members.
(a) Agent reward
(b) Agent reward
(c) Agent penalty
(d) Agent penalty
Figure 9: Agent reward and average team reward during training for (a) fully collaborative scenario and (b) fully competitive scenario. And agent penalty during training for (c) fully collaborative scenario and (d) fully competitive scenario. Each episode represents a full simulation run of the $CO_{2}$ storage over the defined time horizon. The episode index on the x-axis reflects the progression of training, where early episodes capture exploratory behavior and later episodes reflect increasingly optimized policies.
@@
(b) NPV in competitive setting
Figure 10: NPV of each company under (a) fully collaborative scenario and (b) fully competitive scenario. The total team NPV from fully collaborative scenario is \$8702.41 Million, while the total NPV from fully competitive case is \$6875.34 Million.
The proposed framework effectively maintains safe injection pressures across all injection stages, ensuring compliance with injection safety requirements. As illustrated in Figure 11, pressures are monitored at injection well locations for each project area under both fully collaborative and fully competitive scenarios. However, because only point-wise constraints are applied, regional compliance with pressure limits is not guaranteed. This limitation is evident in figure 11 (b), where at the end of injection, the maximum pressure in Company B’s area exceeds its threshold of 55,000 kPa due to pressure migration from neighboring regions, particularly from Company A’s area—a phenomenon referred to as pressure trespass (de Figueiredo, 2005). These highlight the necessity of a system-wide penalty model that accounts for interactions across project boundaries, especially in cases where companies operate under different pressure thresholds and risk tolerances.
(a) Pressure map at selected injection periods, in kPa
(b) Pressure map at selected injection periods, in kPa
Figure 11: Pressure maps at different injection stages t = 5, 10, 20, 25 years (from top to bottom) for (a) fully collaborative scenario and (b) fully competitive scenario. The maximum pressure at each leased project area is annotated.
@@ 5.3 Comparison against MOO results
In this subsection, we assess the performance of the proposed CMG framework relative to conventional MOO algorithms. The Pareto front generated by CMOO for the competitive case is illustrated in figure 12, while Table 2 summarizes the NPV results across different coalition structures and methods. In the fully collaborative scenario, safe MADDPG yields the highest total NPV of \$8720.41 M, exceeding MOO’s best collaborative result (\$8544.07 M) by a notable margin. In the fully competitive setting, MADDPG attains a NPV of \$6875.34 M, which is lower than any of MOO’s competitive allocations, this can be explained that MADDPG in a competitive setting becomes not Pareto optimal due to decentralized structure.
From a per-company perspective, MADDPG provides relatively balanced gains without overly favoring one company at the expense of others, while MOO’s preference-based allocations (favoring A, B, or C) can result in disproportionate benefits to a single entity. Although MADDPG does not guarantee Pareto-optimality for all scenarios, its adaptive learning strategies effectively navigate the trade-off between cooperative and competitive objectives. This makes it a robust choice for maximizing system-wide returns while mitigating excessive inequality among stakeholders.
Figure 12: Pareto front obtained using NSGA-II for the fully competitive case. The units are in \$ million.
@@
MADDPG, fully collab 2451.25 2841.39 3409.77 8720.41
MADDPG, fully comp 2148.92 2985.56 1740.86 6875.34
MOO, fully collab 2039.10 3395.51 3109.46 8544.07
MOO, fully comp(knee point) 1482.99 2989.91 3587.52 8060.41
MOO, fully comp(favoring A) 2069.77 2799.95 2661.47 7531.19
MOO, fully comp(favoring B) 1251.45 3654.18 2216.09 7121.72
MOO, fully comp(favoring C) 1544.31 2319.41 3827.17 7690.90
5.4 Region-wide Penalty Model
In this subsection, we assess the effectiveness of implementing a region-wide penalty, as formulated in Equation 19, to ensure that safety constraints are satisfied across the entire project area. Unlike the point-wise penalty approach, which focuses only on injection well locations, the region-wide penalty enforces compliance at all spatial locations.
Figure 13 illustrates the evolution of reward and penalty during training. Compared to the point-wise penalty case, the region-wide approach yields a denser penalty profile early in training. Over time, penalties steadily decline and stabilize below the threshold, demonstrating that the learned policies achieve safe operation consistently across all regions.
The corresponding pressure maps in Figure 14 confirm that, under both cooperative and competitive scenarios, maximum pore pressures at all injection stages remain below their respective thresholds in every subregion. This highlights not only the robustness of the safe–MADDPG framework but also the value of the region-wide penalty in guiding policy learning toward globally feasible and safe injection strategies, even when multiple agents with potentially conflicting objectives are involved.
(a) Agent reward when region-wide constraints are imposed
(b) Agent reward when region-wide constraints are imposed
(c) Agent penalty when region-wide constraints are imposed
(d) Agent penalty when region-wide constraints are imposed
Figure 13: Agent reward and average team reward during training when region-wide constraints are imposed for (a) fully collaborative scenario and (b) fully competitive scenario. And agent penalty during training for (c) fully collaborative scenario and (d) fully competitive scenario.
(a) Pressure map at selected injection periods when region-wide constraints are imposed, in kPa
(b) Pressure map at selected injection periods when region-wide constraints are imposed, in kPa
Figure 14: Pressure maps at different injection stages t = 5, 10, 20, 25 years (from top to bottom) for (a) fully collaborative scenario and (b) fully competitive scenario. The maximum pressure at each leased project area is annotated. It can be observed that the proposed framework guarantees safe injection at all CCS stages as the maximum pressure at each leased area is less than each company’s threshold pressure as described in 5.1.

### v2

@@ 4.3 Reward and Penalty Formulation for Geological CO2 Storage
To address these challenges, we adopt coalition structures inspired by game theory, enabling flexible combinations of cooperative and competitive behaviors. Coalition structures are enumerated using Bell numbers, allowing evaluation of all possible collaborative configurations. In our case study with three operators (A, B, and C), 5 different coalition structures are considered: the coalition types, structures, and corresponding objectives and reward formulations are summarized in Table 1. The fully cooperative one is where all companies act collectively to maximize the total net present value (NPV) of the project, and fully competitive one is where each company independently seeks to maximize its own NPV without regard for the others. This framework allows us to capture a wide spectrum of strategic behaviors, from joint optimization for mutual benefit to self-interested competition. Based on these coalition types, distinct reward and penalty structures are defined to reflect the corresponding operational incentives and constraints.
Table 1: Coalition types, structures, objectives, and corresponding reward formulations.
Coalition Types Coalition Structure Objective Reward Definition
Fully competitive {A},{B},{C} Each agent $i$ acts independently and maximizes its own financial return. $r^{A}=NPV^{A},\;r^{B}=NPV^{B},\;r^{C}=NPV^{C}$
Fully cooperative/collaborative {A & B & C} Agents A, B, and C form a grand coalition to maximize the collective financial outcome of the team. $r^{A}=r^{B}=r^{C}=NPV^{A}+NPV^{B}+NPV^{C}$
Partially cooperative: coalition AB vs. C {A & B}, {C} Agents A and B cooperate to maximize their joint return, while agent C acts independently to maximize its own NPV. $r^{A}=r^{B}=NPV^{A}+NPV^{B},\;r^{C}=NPV^{C}$
Partially cooperative: coalition AC vs. B {A & C}, {B} Agents A and C cooperate to maximize their coalition reward, while agent B maximizes its own individual return. $r^{A}=r^{C}=NPV^{A}+NPV^{C},\;r^{B}=NPV^{B}$
Partially cooperative: coalition BC vs. A {B & C}, {A} Agents B and C act as a coalition to improve their combined NPV, whereas agent A optimizes individually. $r^{B}=r^{C}=NPV^{B}+NPV^{C},\;r^{A}=NPV^{A}$
For the penalty formulation, the incurred cost is defined to be proportional to the amount by which the pore pressure in the injection-well grid block exceeds a prescribed threshold:
$c^{i}_{t}=5000\cdot\sum_{g\in\mathcal{G}_{well}^{i}}\mathbf{1}\!\left(p_{g}^{t}>p_{thr}\right),$ (21)
@@
In this formulation, $c^{i}_{t}$ denotes the penalty/cost incurred by agent $i$ at time step $t$. The set $\mathcal{G}^{i}$ represents all grid blocks within company $i$’s project area, and $p_{g}^{t}$ is the pressure in grid block $g$ at time $t$. The indicator function $\mathbf{1}(p_{g}^{t}>=p^{i}_{\text{thr}})$ returns $1$ if the pressure in injection-well grid block $g$ exceeds the threshold pressure $p^{i}_{\text{thr}}$ assigned to agent $i$, and 0 otherwise. Thus, the summation counts the total number of violating injection-well grid blocks, which is then multiplied by a fixed penalty factor of $5000$ (5000 is selected in a way that the penalty value is of same magnitude of the reward in order for good training practice) to determine the agent’s cost at that time step.
@@
$\sum_{t=1}^{T}\gamma_{t}\cdot c^{i}(t)\leq d^{i}$
$d^{i}$ is the penalty threshold, in this experiment, because the pressure constraints are strictly imposed, $d^{i}$ is set to be 0.0.
@@ 5.1 Description of Reservoir and Model Setup
The proposed framework is implemented in a reservoir where the dimension spans $44\mathrm{km}\times 12.8\mathrm{km}\times 0.2\mathrm{km}$ in x, y and z directions. With a depth of 4,500m4{,}500\ $\mathrm${m} and an assumed fracture gradient of 17.0kPa/m17.0\ $\mathrm${kPa/m}, the fracture pressure $p_{\mathrm{frac}}$ is set to be 76,500kPa76{,}500\ $\mathrm${kPa}. For safe CO2 injection, the pore pressure must remain below the fracturing pressure, expressed as:
@@
where $0.0<\alpha<1.0$ reflects site-specific risk tolerance. In this study, $\alpha$ should also account for the prediction error between the reduced-order proxy model and the high-fidelity model. The permeability distribution for the basin is shown in Figure 6.
We consider three companies (A, B, and C) operating in non-overlapping regions of the basin (as indicated in the shaded area in Figure 6), each managing a different number of injection wells. To reflect varying risk tolerances, the maximum allowable pore pressure for each company is set to:
$p_{\mathrm{threshold},i}=\alpha_{i}\,p_{\mathrm{frac}}=\begin{cases}75{,}000\ \mathrm{kPa},&i=A,\\ 65{,}000\ \mathrm{kPa},&i=B,\\ 65{,}000\ \mathrm{kPa},&i=C.\end{cases}$
@@
Economic assumptions for revenue and cost calculations are summarized in Table 2, with all companies sharing the same unit revenue and cost structure. Injection rate of each company is varied every 1 year and the total injection lasts 20 years.
@@
5.2 Fully Cooperative vs. Fully Competitive
In this subsection, we present two baseline scenarios: a fully cooperative case in which all companies form a grand coalition to maximize total team NPV, and a fully competitive case where each company operate independently without any coalition (fully competitive case). Figure 7 displays the reward and penalty trajectories over the training process.
In the fully cooperative case, all companies share the same team reward, see Table 1. As observed from Figure 7 (a), it starts with conservative injection operations and improves steadily up to approximately episode 1000, after which the reward curve levels off, indicating convergence towards optimal policies. During early training, the policy may temporarily violate the safety threshold (as reflected by spikes in the penalty curve 7 (c)), but it ultimately stabilizes at a solution that satisfies the safety requirement, demonstrating constraint-compliant operation.
In the fully competitive case, each company also starts conservatively and rapidly improves its own policy during the first 50 episodes. Between roughly episodes 50 and 750, companies adapt to one another’s actions and competitive responses intensify, which reduces the total reward. From about episode 750 to 1500, the policies adjust toward a more balanced outcome and the total reward recovers, eventually stabilizing after episode 1500, consistent with convergence to an equilibrium-like behavior.
Figure 8 summarizes the resulting NPVs under these two settings. The total team NPV increases from \$8942 million in the competitive setting to \$12191 million in the cooperative setting. At the company level, all three companies gain substantial benefits from coalition formation, with NPV gains of \$609 million, \$1255 million and \$1385 million respectively, compared to acting alone. To that end, there are clear incentives for all operators to participate in the grand coalition and to coordinate the operations in the basin.
(a) Per-agent and total reward in fully collaborative scenario
(b) Per-agent and total reward in fully competitive scenario
(c) Per-agent penalty in fully collaborative scenario
(d) Per-agent penalty in fully competitive scenario
Figure 7: Per-agent and total reward during training for (a) the fully cooperative scenario and (b) the fully competitive scenario, along with the corresponding per-agent penalties for (c) the fully cooperative scenario and (d) the fully competitive scenario. Each episode represents a complete simulation run of the CO2 storage over the defined time horizon. The episode index reflects training progress: early episodes exhibit more exploratory behavior, whereas later episodes correspond to increasingly optimized policies.
@@
(b) NPVs in competitive setting
Figure 8: NPV of individual company under (a) fully collaborative scenario and (b) fully competitive scenario. The total team NPV is is \$12191 Million for the fully cooperative case and \$8942 Million for the fully competitive case..
Overall, the proposed framework maintains injection pressures within safe limits throughout the injection period, ensuring compliance with operational safety requirements.. As shown in Figure 9, the maximum pressures at injection-well locations remain below the fracturing pressure at the end of injection in both the fully cooperative and fully competitive settings. Notably, grand coordination enables substantially better utilization of the available pressure space, particularly for Company B, highlighting the effectiveness of the proposed approach. In contrast, under full competition, self-interested responses to opponents’ actions prevent efficient use of the available pressure space.
(a) Pressure map (fully cooperative) at selected injection periods, in kPa
(b) Pressure map (fully competitive) at selected injection periods, in kPa
Figure 9: Pressure maps at different injection stages t = 5, 10, 15, 20 years (from top to bottom) for (a) fully collaborative scenario and (b) fully competitive scenario. The maximum pressure reached at each leased project area is annotated.
5.3 Mixed Cooperative and Competitive
In this subsection, we examine partially collaborative settings in which some companies cooperate while competing with the remaining ones. This leads to three coalition structures: (i) A and B cooperate while competing with C, (ii) A and C cooperate while competing with B, and (iii) B and C cooperate while competing with A. Figure 10 summarizes the corresponding reward trajectories during training.
Across all three mixed-coalition scenarios, a consistent pattern emerges that echoes the previous subsection: forming a coalition provides a systematic advantage over operating alone. When Companies A and B form a coalition against C, the coalition members quickly converge to a stable, relatively high reward, whereas the independent agent C settles into a lower but steady payoff as it adapts to the coalition’s coordinated behavior. A similar dynamic appears when A cooperates with C against B: the coalition partners achieve high long-term rewards, whereas the solo competitor B experiences a gradual decline in performance as the coalition’s joint strategy strengthens. Finally, when B and C cooperate against A, the coalition again maintains a sustained reward advantage, and A converges to a lower equilibrium level after an initial transient. Overall, these results demonstrate that cooperation—even when partial—can create persistent performance gaps, with coalition members consistently outperforming the isolated competitor in the long-run equilibrium.
(a) Per-agent and total reward in mixed {A, B} and {C} scenario
(b) Per-agent and total reward in mixed {A, C} and {B} scenario
(c) Per-agent and total reward in mixed {B, C} and {A} scenario
Figure 10: Per-agent and total reward trajectories during training for the three mixed-coalition settings: (a) {A, B} v.s {C}, (b) {A, C} v.s {B} and (c) {B, C} v.s {A}.
@@ 5.4 Comparison against MOO results
In this subsection, we assess the performance of the proposed constrained Markov game (CMG) framework benchmarked by constrained MOO solutions. The Pareto front produced by NSGA-II is shown in Figure 11, and Table 3 summarizes the NPVs obtained under different coalition structures and methods.
As suggested in Table 3, the fully cooperative MADDPG setting yields the highest NPVs for all three companies (A: \$2367 M, B: \$4876 M, C: \$4948 M) among different coalition structures. The total NPV reaches its maximum (\$12,191 M) in this setting, showing that end-to-end policy learning can discover highly coordinated strategies that improve overall system efficiency when operators share a common objective. In contrast, fully competitive MADDPG results in substantial decrease of payoffs for each company—most notably for Company C (a \$1385 M decrease compared to the grand coalition setting). The grand competition case also leads to the lowest overall NPV (\$8942 M), illustrating the inefficiency induced when agents optimize selfish objectives without coordination.
The mixed-coalition MADDPG settings create asymmetric value distributions: when A & B cooperate against C, both coalition members improve relative to the competitive baseline, especially B (\$4563 M), while C receives similar reward relative to the grand competitive case. Similarly, in A & C vs. B, the coalition partners (A and C) gain moderately, whereas B is disadvantaged (\$3597 M). In B & C vs. A, coalition members B and C capture very high NPVs (4750 and 4766), while A receives a low payoff (\$1808 M). Collectively, the mixed-coalition outcomes highlight a consistent distributional effect: forming a coalition increases the coalition members’ NPVs relative to the fully competitive baseline, while the remaining standalone operator is comparatively disadvantaged. In particular, the paired firms in each mixed setting capture most of the value created by coordinated actions, whereas the non-coalition company converges to a lower equilibrium payoff. These results suggest that partial cooperation can yield substantial coalition gains, but it may also amplify payoff asymmetries across operators under decentralized competition.
By comparison, the MOO baseline generates a set of non-dominated trade-off solutions. We note, however, that NSGA-II is a heuristic and may not recover the exact Pareto set in all cases [19, 54]; here it is therefore used as a practical benchmark. The knee point selection yields relatively high NPVs for all three companies (A: \$2419 M, B: \$4630 M, C: \$4415 M) and a strong total NPV (\$11 464 M). When the selection criterion is adjusted to favor a single company, the favored firm’s NPV increases accordingly, as reflected in the A-favored (\$2475 M), B-favored (\$4709 M), and C-favored (\$4618 M) cases.
Although the cooperative MARL setting optimizes a team reward while MOO explicitly parameterizes multi-objective trade-offs, Table 3 indicates that the fully cooperative MADDPG outcome is non-dominated within the reported set: it achieves the largest total NPV and attains the highest individual NPVs for Companies B and C (while Company A is slightly lower than the MOO knee-point result). Therefore, within the discrete set of solutions compared in Table 3, the cooperative MADDPG result is consistent with a Pareto-efficient outcome. We emphasize that this observation does not constitute a formal proof of global Pareto optimality over the continuous feasible space, which would require verifying dominance against the full Pareto front.
Overall, MADDPG is effective at discovering highly coordinated policies that can maximize coalition-level gains under different coalition structures, whereas MOO provides interpretable, preference-controlled trade-offs that may sacrifice some peak system performance in exchange for more predictable and stakeholder-balanced allocations.
Figure 11: Pareto front obtained using NSGA-II for constrained MOO. The units are in \$ million.
@@
MADDPG, fully collab 2367 4876 4948 12191
MADDPG, fully comp 1758 3621 3563 8942
MADDPG, A&B v.s. C 2316 4563 3624 10503
MADDPG, A&C v.s. B 2300 3597 4557 10454
MADDPG, B&C v.s. A 1808 4750 4766 11324
MOO, knee point 2419 4630 4415 11464
MOO, favoring A 2475 4295 4238 11008
MOO, favoring B 2370 4709 4180 11259
MOO, favoring C 2106 4404 4618 11128